OpenStudy (anonymous):

Find the derivative of the function: h(x)= 6^x^2-x

OpenStudy (camper4834):

\[6^{x ^{2}-x}\] or \[6^{x ^{2}}-x\]

OpenStudy (anonymous):

Its the first one

OpenStudy (camper4834):

so I'm thinking of starting out (6^x^2)/(6^x)

OpenStudy (camper4834):

then that becomes \[\frac{ e ^{\ln(6)x ^{2}} }{ e ^{\ln(6)x} }\] then do low D hi - hi D low over the sqaure of whats below

OpenStudy (anonymous):

Ok great thank you so I'm following your work so far, what comes next after that?

OpenStudy (camper4834):

Just long work \[\frac{ 6^{x}*2x*\ln(6)6^{x ^{2}} - 6^{x ^{2}}*\ln(6)6^{x} }{ (6^{x})^{2} }\] \[=\frac{ \ln(6)6^{x ^{2}}(2x-1) }{ 6^{x} }\] this is the cleanest i got it

OpenStudy (anonymous):

oh ok looks good I see what you did there on top